# Unraveling the Smoothness Properties of Diffusion Models: A Gaussian Mixture Perspective

### v1

@@ F.1 Definitions
F.2 Lemmas for calculation of the score function
@@
where the first step follows from Definition F.1, the second step follows from Fact C.1, the third step follows from Fact C.3, and the last step follows from simple algebra. ∎
@@
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition F.1.
@@
where the first step follows from Definition F.2, the second step follows from Fact C.1, and the last step follows from Lemma F.3. ∎
F.3 Calculation of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition F.1.
  • Let $f(x),g(x)$ be defined as Definition F.5.
@@
where the first step follows from Fact C.1, the second step follows from Fact C.1, the third step follows from Lemma F.4, the fourth step follows from Definition F.2 and the last step follows from Definition F.5.
F.4 Lemmas for the calculation of the upper bound of the score function
@@
F.5 Upper bound of the score function
@@
Lemma F.8 (Linear growth).
@@
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $f(x),g(x)$ be defined as Definition F.5.
@@
where the first step follows from Lemma F.6, the second step follows from Fact C.2, the third step follows from $f(x)+g(x)=1$ and $f(x),g(x)\geq 0$, the fourth step follows from Lemma F.7, the fifth step follows from definition of $\mu_{\max}$ and $\sigma_{\min}$, and the last step follows from $\mu_{\max}\geq 1$. ∎
F.6 Lemmas for Lipschitz calculation: upper bound of base functions
@@
where the first step follows from Fact C.2, the second step follows from Lemma F.10. ∎
F.7 Lemmas for Lipschitz calculation: Lipschitz constant of base functions
@@
where the first step follows from Fact C.2, the second step follows from simple algebra, the third step follows from Fact C.2, the fourth step follows from $\|x-\mu_{i}(t)\|_{2}\leq R$, $\|\widetilde{x}-\mu_{i}(t)\|_{2}\leq R$, the fifth step follows from Lemma F.12, and the last step follows from simple algebra. ∎
@@
where the first step follows from Definition F.1, the second step follows from simple algebra, the third step follows from Fact C.9, and the last step follows from Lemma F.11. ∎
@@
where the first step follows from simple algebra, the second step follows from Fact C.2, the third step follows from Lemma F.14, the fourth step follows from $\alpha\in(0,1)$, and the last step follows from the definition of $\det_{\min},\sigma_{\max},\sigma_{\min}$. ∎
@@
where the first step follows from simple algebra, the second step follows from $\alpha N_{1}(x)+(1-\alpha)N_{2}(x)\geq\gamma$, and the last step follows from Lemma F.15. ∎
F.8 Lemmas for Lipschitz calculation: $f(x)$
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition F.1.
  • Let $f(x)$ be defined as Definition F.5.
@@
where the first step follows from $N_{1}(x)\leq\frac{1}{(2\pi)^{d/2}\det(\Sigma_{1}(t))^{1/2}}$, the second step follows from Lemma F.16 and the last step follows from definition of $\det_{\min}$.
@@
where the first step follows from $\alpha N_{1}(x)+(1-\alpha)N_{2}(x)\geq\gamma$, the second step follows from Lemma F.15.
Combining Eq. (F.8) and Eq. (F.8) together, we have
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition F.1.
  • Let $f(x)$ be defined as Definition F.5.
@@
where the first step follows from $f(x)\leq 1$, the second step follows from Lemma F.12.
@@
where the first step follows from Lemma F.9, the second step follows from Lemma F.17.
Combining Eq. (F.8) and Eq. (F.8) together, we have
@@
F.9 Lemmas for Lipschitz calculation: $g(x)$
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition F.1.
  • Let $g(x)$ be defined as Definition F.5.
@@
where the first step follows from $N_{2}(x)\leq\frac{1}{(2\pi)^{d/2}\det(\Sigma_{2}(t))^{1/2}}$, the second step follows from Lemma F.16.
@@
where the first step follows from $\alpha N_{1}(x)+(1-\alpha)N_{2}(x)\geq\gamma$, the second step follows from Lemma F.15.
Combining Eq. (F.9) and Eq. (F.9) together, we have
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition F.1.
  • Let $g(x)$ be defined as Definition F.5.
@@
where the first step follows from $g(x)\leq 1$, the second step follows from Lemma F.12.
@@
where the first step follows from Lemma F.9, the second step follows from Lemma F.19.
Combining Eq. (F.9) and Eq. (F.9) together, we have
@@
F.10 Lipschitz constant of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $f(x),g(x)$ be defined as Definition F.5.
@@
where the first step follows from Lemma F.6, the second step follows from Fact C.2, the third step follows from Lemma F.18, the fourth step follows from Lemma F.20, and the last step follows from simple algebra. ∎
Appendix G A General Version for $k$ Gaussian
@@
  • Section G.1 provides the definition for $k$ mixture of Gaussians.
  • Section G.2 provides the expression of the score function.
  • Section G.3 provides the upper bound of the score function.
  • Section G.4 provides lemmas that are used in further calculation of Lipschitz constant.
  • Section G.5 provides the Lipschitz constant for $k$ mixture of Gaussians.
G.1 Definitions
@@
  • Let $N_{i}(x)$ be defined as Definition G.1.
  • Let $p_{t}(x)$ be defined as Definition G.2
@@
where the first step follows from Definition G.2, the second step follows from Fact C.1, and the last step follows from Lemma F.3.
@@
G.2 Calculation of the score function
@@
  • Let $N_{i}(x)$ be defined as Definition G.1.
  • Let $p_{t}(x)$ be defined as Definition G.2.
  • Let $f_{i}(x)$ be defined as Definition G.4.
@@
where the first step follows from Fact C.1, the second step follows from Fact C.1, the third step follows from Lemma G.3, the fourth step follows from Definition G.2, and the last step follows from Definition G.4. ∎
G.3 Upper bound of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition G.2.
  • Let $f_{i}(x)$ be defined as Definition G.4.
@@
where the first step follows from Lemma G.5, the second step follows from triangle inequality, the third step follows from $\sum_{i=1}^{k}f_{i}(x)=1$ and $f_{i}(x)\geq 0$, the fourth step follows from Lemma F.7, the fifth step follows from definition of $\mu_{\max}$ and $\sigma_{\min}$, and the last step follows from $\mu_{\max}\geq 1$. ∎
G.4 Lemmas for Lipshitz calculation
@@
where the first step follows from simple algebra, the second step follows from Fact C.2, the third step follows from Lemma F.14, the fourth step follows from $\sum_{i=1}^{k}\alpha_{i}(t)=1$, and $\alpha_{i}(t)\in(0,1)$, and the last step follows from the definition of $\det_{\min},\sigma_{\max},\sigma_{\min}$. ∎
@@
where the first step follows from simple algebra, the second step follows from $\sum_{i=1}^{k}\alpha_{i}(t)N_{i}(x)\geq\gamma$, and the last step follows from Lemma G.7. ∎
@@
  • Let $N_{i}(x)$ be defined as Definition G.1.
  • Let $f_{i}(x)$ be defined as Definition G.4.
@@
where the first step follows from Definition G.4, the second step follows from Fact C.2, and the last step follows from simple algebra.
@@
where the first step follows from $N_{i}(x)\leq\frac{1}{(2\pi)^{d/2}\det(\Sigma_{i}(t))^{1/2}}$, the second step follows from Lemma G.8, and the last step follows from definition of $\det_{\min}$.
@@
where the first step follows from $\sum_{i=1}^{k}\alpha_{i}(t)N_{i}(x)\geq\gamma$, the second step follows from Lemma G.7.
Combining Eq. (G.4) and Eq. (G.4) together, we have
@@
  • Let $N_{i}(x)$ be defined as Definition G.1.
  • Let $f_{i}(x)$ be defined as Definition G.4.
@@
where the first step follows from Lemma F.12.
@@
where the first step follows from Lemma F.9, the second step follows from Lemma G.9.
Combining Eq. (G.4) and Eq. (G.4) together, we have
@@
G.5 Lipschitz constant of the score function
@@
  • Let $N_{i}(x)$ be defined as Definition G.1.
  • Let $p_{t}(x)$ be defined as Definition G.2.
  • Let $f_{i}(x)$ be defined as Definition G.4.
@@
where the first step follows from Lemma G.5, the second step follows from triangle inequality, the third step follows from Lemma G.10, and the last step follows from $\sum_{i=1}^{k}|f_{i}(x)|=\sum_{i=1}^{k}f_{i}(x)=1$ and $\sum_{i=1}^{k}\alpha_{i}(t)=1$. ∎
Appendix H Putting It All Together
Our overall goal is that we want to provide a more concrete calculation for theorems in Section D by assuming the data distribution is a $k$ mixture of Gaussian. Now we provide lemmas that are used in further calculation.
@@ Appendix H Putting It All Together
Lemma H.1 (Formal version of Lemma 3.2).
@@
Lemma H.2 (Formal version of Lemma 3.5).
@@
Lemma H.3 (Formal version of Lemma 3.3).
@@
Using Lemma G.11 and H.1, we can get the result. ∎
Appendix I More Calculation for Application
In this section, we will provide a more concrete calculation for Theorem D.5, Theorem D.6, Theorem D.7, Theorem D.8 and Theorem D.9.
I.1 Concrete calculation of Theorem D.5
Theorem I.1 ($\mathsf{DDPM}$, total variation, formal version of Theorem 5.6).

### v2

@@ E.1 Definitions
E.2 Lemmas for calculation of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition E.2.
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition E.1.
@@
where the first step follows from Definition E.2, the second step follows from Fact C.1, and the last step follows from Lemma E.3. ∎
E.3 Calculation of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition E.2.
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition E.1.
  • Let $f(x),g(x)$ be defined as Definition E.5.
@@
where the first step follows from Fact C.1, the second step follows from Fact C.1, the third step follows from Lemma E.4, the fourth step follows from Definition E.2 and the last step follows from Definition E.5.
E.4 Lemmas for the calculation of the upper bound of the score function
@@
E.5 Upper bound of the score function
@@
Lemma E.8 (Linear growth).
@@
  • Let $p_{t}(x)$ be defined as Definition E.2.
  • Let $f(x),g(x)$ be defined as Definition E.5.
@@
where the first step follows from Lemma E.6, the second step follows from Fact C.2, the third step follows from $f(x)+g(x)=1$ and $f(x),g(x)\geq 0$, the fourth step follows from Lemma E.7, the fifth step follows from definition of $\mu_{\max}$ and $\sigma_{\min}$, and the last step follows from $\mu_{\max}\geq 1$. ∎
E.6 Lemmas for Lipschitz calculation: upper bound of base functions
@@
where the first step follows from Fact C.2, the second step follows from Lemma E.10. ∎
E.7 Lemmas for Lipschitz calculation: Lipschitz constant of base functions
@@
where the first step follows from Fact C.2, the second step follows from simple algebra, the third step follows from Fact C.2, the fourth step follows from $\|x-\mu_{i}(t)\|_{2}\leq R$, $\|\widetilde{x}-\mu_{i}(t)\|_{2}\leq R$, the fifth step follows from Lemma E.12, and the last step follows from simple algebra. ∎
@@
where the first step follows from Definition E.1, the second step follows from simple algebra, the third step follows from Fact C.9, and the last step follows from Lemma E.11. ∎
@@
where the first step follows from simple algebra, the second step follows from Fact C.2, the third step follows from Lemma E.14, the fourth step follows from $\alpha\in(0,1)$, and the last step follows from the definition of $\det_{\min},\sigma_{\max},\sigma_{\min}$. ∎
@@
where the first step follows from simple algebra, the second step follows from $\alpha N_{1}(x)+(1-\alpha)N_{2}(x)\geq\gamma$, and the last step follows from Lemma E.15. ∎
E.8 Lemmas for Lipschitz calculation: $f(x)$
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition E.1.
  • Let $f(x)$ be defined as Definition E.5.
@@
where the first step follows from Definition E.5, the second step follows from Fact C.2, and the last step follows from simple algebra.
@@
where the first step follows from $N_{1}(x)\leq\frac{1}{(2\pi)^{d/2}\det(\Sigma_{1}(t))^{1/2}}$, the second step follows from Lemma E.16 and the last step follows from definition of $\det_{\min}$.
@@
where the first step follows from $\alpha N_{1}(x)+(1-\alpha)N_{2}(x)\geq\gamma$, the second step follows from Lemma E.15.
Combining Eq. (E.8) and Eq. (E.8) together, we have
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition E.1.
  • Let $f(x)$ be defined as Definition E.5.
@@
where the first step follows from $f(x)\leq 1$, the second step follows from Lemma E.12.
@@
where the first step follows from Lemma E.9, the second step follows from Lemma E.17.
Combining Eq. (E.8) and Eq. (E.8) together, we have
@@
E.9 Lemmas for Lipschitz calculation: $g(x)$
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition E.1.
  • Let $g(x)$ be defined as Definition E.5.
@@
where the first step follows from Definition E.5, the second step follows from Fact C.2, and the last step follows from simple algebra.
@@
where the first step follows from $N_{2}(x)\leq\frac{1}{(2\pi)^{d/2}\det(\Sigma_{2}(t))^{1/2}}$, the second step follows from Lemma E.16.
@@
where the first step follows from $\alpha N_{1}(x)+(1-\alpha)N_{2}(x)\geq\gamma$, the second step follows from Lemma E.15.
Combining Eq. (E.9) and Eq. (E.9) together, we have
@@
  • Let $N_{1}(x),N_{2}(x)$ be defined as Definition E.1.
  • Let $g(x)$ be defined as Definition E.5.
@@
where the first step follows from $g(x)\leq 1$, the second step follows from Lemma E.12.
@@
where the first step follows from Lemma E.9, the second step follows from Lemma E.19.
Combining Eq. (E.9) and Eq. (E.9) together, we have
@@
E.10 Lipschitz constant of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition E.2.
  • Let $f(x),g(x)$ be defined as Definition E.5.
@@
where the first step follows from Lemma E.6, the second step follows from Fact C.2, the third step follows from Lemma E.18, the fourth step follows from Lemma E.20, and the last step follows from simple algebra. ∎
Appendix F A General Version for $k$ Gaussian
@@
  • Section F.1 provides the definition for $k$ mixture of Gaussians.
  • Section F.2 provides the expression of the score function.
  • Section F.3 provides the upper bound of the score function.
  • Section F.4 provides lemmas that are used in further calculation of Lipschitz constant.
  • Section F.5 provides the Lipschitz constant for $k$ mixture of Gaussians.
F.1 Definitions
@@
  • Let $N_{i}(x)$ be defined as Definition F.1.
  • Let $p_{t}(x)$ be defined as Definition F.2
@@
where the first step follows from Definition F.2, the second step follows from Fact C.1, and the last step follows from Lemma E.3.
@@
F.2 Calculation of the score function
@@
  • Let $N_{i}(x)$ be defined as Definition F.1.
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $f_{i}(x)$ be defined as Definition F.4.
@@
where the first step follows from Fact C.1, the second step follows from Fact C.1, the third step follows from Lemma F.3, the fourth step follows from Definition F.2, and the last step follows from Definition F.4. ∎
F.3 Upper bound of the score function
@@
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $f_{i}(x)$ be defined as Definition F.4.
@@
where the first step follows from Lemma F.5, the second step follows from triangle inequality, the third step follows from $\sum_{i=1}^{k}f_{i}(x)=1$ and $f_{i}(x)\geq 0$, the fourth step follows from Lemma E.7, the fifth step follows from definition of $\mu_{\max}$ and $\sigma_{\min}$, and the last step follows from $\mu_{\max}\geq 1$. ∎
F.4 Lemmas for Lipshitz calculation
@@
where the first step follows from simple algebra, the second step follows from Fact C.2, the third step follows from Lemma E.14, the fourth step follows from $\sum_{i=1}^{k}\alpha_{i}(t)=1$, and $\alpha_{i}(t)\in(0,1)$, and the last step follows from the definition of $\det_{\min},\sigma_{\max},\sigma_{\min}$. ∎
@@
where the first step follows from simple algebra, the second step follows from $\sum_{i=1}^{k}\alpha_{i}(t)N_{i}(x)\geq\gamma$, and the last step follows from Lemma F.7. ∎
@@
  • Let $N_{i}(x)$ be defined as Definition F.1.
  • Let $f_{i}(x)$ be defined as Definition F.4.
@@
where the first step follows from $N_{i}(x)\leq\frac{1}{(2\pi)^{d/2}\det(\Sigma_{i}(t))^{1/2}}$, the second step follows from Lemma F.8, and the last step follows from definition of $\det_{\min}$.
@@
where the first step follows from $\sum_{i=1}^{k}\alpha_{i}(t)N_{i}(x)\geq\gamma$, the second step follows from Lemma F.7.
Combining Eq. (F.4) and Eq. (F.4) together, we have
@@
  • Let $N_{i}(x)$ be defined as Definition F.1.
  • Let $f_{i}(x)$ be defined as Definition F.4.
@@
where the first step follows from Lemma E.12.
@@
where the first step follows from Lemma E.9, the second step follows from Lemma F.9.
Combining Eq. (F.4) and Eq. (F.4) together, we have
@@
F.5 Lipschitz constant of the score function
@@
  • Let $N_{i}(x)$ be defined as Definition F.1.
  • Let $p_{t}(x)$ be defined as Definition F.2.
  • Let $f_{i}(x)$ be defined as Definition F.4.
@@
where the first step follows from Lemma F.5, the second step follows from triangle inequality, the third step follows from Lemma F.10, and the last step follows from $\sum_{i=1}^{k}|f_{i}(x)|=\sum_{i=1}^{k}f_{i}(x)=1$ and $\sum_{i=1}^{k}\alpha_{i}(t)=1$. ∎
Appendix G Putting It All Together
Our overall goal is that we want to provide a more concrete calculation for theorems in Section 6 by assuming the data distribution is a $k$ mixture of Gaussian. Now we provide lemmas that are used in further calculation.
@@ Appendix G Putting It All Together
Lemma G.1 (Formal version of Lemma 3.2).
@@
Lemma G.2 (Formal version of Lemma 3.5).
@@
Lemma G.3 (Formal version of Lemma 3.3).
@@
Using Lemma F.11 and G.1, we can get the result. ∎
Appendix H More Calculation for Application
In this section, we will provide a more concrete calculation for Theorem 6.5, Theorem 6.6, Theorem 6.7, Theorem 6.8 and Theorem 6.9.
H.1 Concrete calculation of Theorem 6.5
Theorem H.1 ($\mathsf{DDPM}$, total variation, formal version of Theorem 5.6).